# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v2

@@ 3.2 Predictive risk
As $\lambda\rightarrow+\infty$, the limit of the deterministic equivalent $r_{\lambda}(\Upsilon_{n})$ of the predictive risk is
$\lim_{\lambda\rightarrow+\infty}r_{\lambda}(\Upsilon_{n})=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}.$
Moreover, as $\lambda\rightarrow 0$, the limit of the deterministic equivalent $r_{\lambda}(\Upsilon_{n})$ is as follows
@@
$\lim_{\lambda\rightarrow 0,\;\lambda>0}r_{\lambda}(\Upsilon_{n})=\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}T_{p}(0^{-})]+\sigma^{2},$
@@
$\lim_{\lambda\rightarrow 0,\;\lambda>0}r_{\lambda}(\Upsilon_{n})=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}\kappa(0^{+})(\Sigma_{n}+\kappa(0^{+}))^{-1}]+\frac{\sigma^{2}}{n}\mathrm{Tr}[\kappa^{\prime}(0^{+})\widetilde{S}_{p}\Sigma_{n}(\Sigma_{n}+\kappa(0^{+}))^{-2}]+\sigma^{2},$
where $\kappa^{\prime}(\lambda)$ denotes the derivative with respecto to $\lambda$ of the diagonal matrix $\kappa(\lambda)$ defined in Equation (3.2).
In the case $\lambda\rightarrow 0$, Corollary 3.2 yields a deterministic equivalent of the predicitve risk of the mininum norm least-square estimator $\hat{\theta}$. Then, it can be seen that the risk $r_{0}(\Upsilon_{n})$ exhibit two different behaviors depending on the value of the ratio $p/n$ with respect to one.
If $p/n<1$, then $\hat{\theta}$ is equivalent to the ordinary least-square estimator $(X_{n}^{t}X_{n})^{-1}X_{n}^{t}Y_{n}$ which is known to be unbiased, and the risk $r_{0}(\Upsilon_{n})$ is thus only composed of a variance term. Moreover, if the variance profile is assumed to be quasi-bistochastic, then, given that
@@
$r_{0}(\Upsilon_{n})=\frac{\sigma^{2}}{1-p/n}\times\frac{1}{n}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}=\sigma^{2}\frac{p/n}{1-p/n}+\sigma^{2}\rightarrow\sigma^{2}\frac{c}{1-c}+\sigma^{2},\;\mbox{as}\;n\to\infty,\;p/n\to c,$
which corresponds, when $c<1$, to the known asymptotic limit of the predictive risk of least squares estimation for iid data when the entries of the features matrix $X_{n}$ are iid centered random variables with variances equal to $1$, see e.g. [HMRT22][Proposition 2].
@@
$r_{0}(\Upsilon_{n})=\alpha^{2}\left(1-\frac{n}{p}\right)+\sigma^{2}\frac{1}{p/n-1}\rightarrow\alpha^{2}\left(1-\frac{1}{c}\right)+\sigma^{2}\frac{1}{c-1},\;\mbox{as}\;n\to\infty,\;p/n\to c,$
which corresponds, when $c>1$, to known results on the bias-variance decomposition of the asymotitic limit of the predictive risk of the minimum norm least squares estimator for iid data when the entries of $X_{n}$ are iid centered random variables with variance $1$, see e.g. [HMRT22][Theorem 1].
Beyond the assumption of a quasi-stochastic variance profile, it is difficult to analytically determine the shape of the preditive risk $r_{0}(\Upsilon_{n})$ as a function of the ratio $p/n$. Indeed, this requires to at least know upper and lower bounds on the magnitude of the diagonal elements of $T_{p}(0^{-})$ and $\kappa(0^{+})$ (and its derivative). As shown by Lemma 3.2 when $p/n<1$, this issue amounts to finding upper and lower bounds of the support of the matrix-valued measure $\mu$ satisfying
$T_{p}(-0)=\int_{\tau}^{+\infty}\frac{\mu(dw)}{w}.$
Hence, upper and lower bounding $T_{p}(-0)$ is related to understanding the value of the constant $\tau$ and the size of the support of the limiting spectral distribution of the covariance matrice $\Sigma_{n}$ which remains (to the best of our knowledge) an open problem for random matrices with an arbitrary variance profile.
@@
The function $\lambda\mapsto\hat{r}_{\lambda}(X_{n})$ reaches its minimium at $\lambda_{\ast}=\frac{\sigma^{2}p}{\alpha^{2}n}$ meaning that
$\hat{r}_{\lambda_{\ast}}(X_{n})\leq\hat{r}_{\lambda}(X_{n}),$ (3.5)
for any $\lambda>0$.
Note that although $\hat{r}_{\lambda}(X_{n})$ depends on the variance profile, this is not the case of the optimal value of $\lambda_{\ast}$ that minimizes the predictive risk as shown by Corollary 3.3. Moreover, the optimal value $\lambda_{\ast}$ is also the one minimizing $\lambda\mapsto\hat{r}_{\lambda}(X_{n})$ in the framework of [DW18] with a matrix $X_{n}$ of features made of iid rows.
@@ 4 Numerical experiments
In this section, we illustrate the results of this paper with numerical experiments. Although we do not have an explicit formula for $T_{p}(z)$, this matrix-valued function satisfies the fixed-point equation (2.3). This allows us to approximate $T_{p}(z)$ numerically using a fixed-point algorithm. In this manner, we are thus able to obtain a numerical approximation of the deterministic equivalent $r_{\lambda}(\Upsilon_{n})$ of the predictive risk. In this section, we consider several variance profiles that we normalize such that
@@
$\Gamma_{n}=\left[\begin{array}[]{ccc}\gamma_{1}^{2}\mathbbm{1}_{n/4}\mathbbm{1}_{p/4}^{t}&\gamma_{2}^{2}\mathbbm{1}_{n/4}\mathbbm{1}_{p/3}^{t}&\mathbbm{1}_{n/4}\mathbbm{1}_{5p/12}^{t}\\ \gamma_{2}^{2}\mathbbm{1}_{n/3}\mathbbm{1}_{p/4}^{t}&\gamma_{1}^{2}\mathbbm{1}_{n/3}\mathbbm{1}_{p/3}^{t}&\gamma_{3}^{2}\mathbbm{1}_{n/3}\mathbbm{1}_{5p/12}^{t}\\ \mathbbm{1}_{5n/12}\mathbbm{1}_{p/4}^{t}&\gamma_{3}^{2}\mathbbm{1}_{5n/12}\mathbbm{1}_{p/3}^{t}&\gamma_{1}^{2}\mathbbm{1}_{5n/12}\mathbbm{1}_{5p/12}^{t}\end{array}\right],$
@@
The values of the predictive risk and its deterministic equivalent are compared for various variance profiles in Figure 3 with $\lambda$ ranging from 0.1 to 10, $n=400$ and $p=600$. The curves displayed in Figure 3 confirm that $r_{\lambda}(\Upsilon_{n})$ is a very accurate estimator of $\hat{r}_{\lambda}(X_{n})$ in high-dimension since the dashed and solid lines coincide. These variance profiles also provide curves of predictive risks having similar shapes (up to a vertical translation). The curves representing the cases of constant and quasi doubly stochastic variance profile coincide, confirming the comment made in Section 1.1 on the doubly stochastic profile and its similarities with the setting of iid data. Moreover, the minimum of $r_{\lambda}(\Upsilon_{n})$ is indeed reached at the optimal value $\lambda_{\ast}$ for any variance profile as shown by Corollary 3.3.
Figure 3: Comparison of $\hat{r}_{\lambda}(X_{n})$ and $r_{\lambda}(\Upsilon_{n})$ for several variance profiles , $\alpha=1$, $\sigma=1$, $n=400$ and $p=600$. The dashed lines correspond to the predictive risk while the solid lines correspond to the deterministic equivalent.
We illustrate the appearance of a double or triple descent phenomenon for several variance profiles in Figure 4 (a), (c) and (e). These figures represent the predictive risk and its approximation by a deterministic equivalent for various values of the ratio $p/n$ with $\lambda=0$. The solid lines depicts $r_{0}(\Upsilon_{n})$ whereas the dashed lines represent $\hat{r}_{0}(X_{n})$. For every variance profile, the solid and dashed lines coincide which confirms that $r_{0}(\Upsilon_{n})$ is a relevant deterministic equivalent of $\hat{r}_{0}(X_{n})$. For the constant and quasi doubly stochastic variance profiles, we observe the well known double descent phenomenon as illustrated by Figure 4 (a) since the curves are increasing for $p/n<1$ and decreasing for $p/n>1$. This is related to Corollary 3.2 which states that the expression of $\lim_{\lambda\rightarrow 0}r_{\lambda}(\Upsilon_{n})$ depends upon the value of the ratio $p/n$ with respect to one. Nevertheless, for other variance profiles, the shape of the predictive risk can be very different from one variance profile to another, and it differs from the usual double descent. For some variance profiles a phenomenon of triple descent arises, notably for the piecewise constant and block cases, as shown in Figure 4 (c) and (e). We also remark the appearance of a quadruple descent in the case of the polynomial profile, as shown in Figure 4 (e).
@@ 5 Conclusion
In this paper, we have derived a deterministic equivalent of the DOF and the predictive risk of ridge (less) regression in a high-dimensional framework with a variance profile to handle the setting of non-iid data. In order to achieve this result, we had to use ingredients from RMT to determine a deterministic equivalent of the diagonal of $Q_{p}(z)$. We have shown that results on the values of the DOF and the predictive that are already known for a constant variance profile, notably those of [DW18] and [Bac24], also hold in the case of a quasi doubly stochastic profile. However, there are variance profiles that cause a behavior of the predictive risk of the minimum norm least square estimator that differs from the double descent phenomenon as classically observed in the case of a constant variance profile. The numerical experiments that we have conducted confirm that our deterministic equivalent accurately estimates the predictive risk in high-dimension. Our results also allow to understand how assuming such a variance profile for the data influences the statistical properties of ridge regression when compared to the standard assumption of iid observations. We hope that our approach on the use of variance profiles may lead to further research works on the statistical analysis of other estimators than ridge regression in more complex models with non-iid data.
@@
A.1 Proofs of Proposition 3.1 and Lemma 3.1
@@
Recall that $\frac{1}{p}\widehat{df}_{1}(\lambda)=\frac{1}{p}\operatorname{Tr}[\widehat{\Sigma}_{n}(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}]$, for $\lambda>0$. Thus by using $\hat{\mu}_{n}$ the empirical eigenvalue distribution of $\widehat{\Sigma}_{n}$ we have this new expression of the degree of freedom
@@
Moreover, $\hat{\mu}_{n}$ being a probability measure one can notice from Definition 2.1 that
@@
Thanks to Proposition 2.1, one has that
$\frac{1}{p}\widehat{df}_{1}(\lambda)\sim 1-\lambda g_{\nu_{n}}(-\lambda).$ (A.1)
Let us now remark that, by Proposition 2.1, one has that
@@
where
$\kappa_{j}(\lambda)=\frac{\operatorname{Tr}[\widetilde{D}_{j}]}{\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(-\lambda)]},\quad\mbox{for}\quad 1\leq j\leq p,$
and
$\kappa(\lambda)=\mathop{\rm diag}\nolimits(\kappa_{1}(\lambda),\ldots,\kappa_{p}(\lambda)).$
Consequently, by combining Equations (A.1) and (A.2), we obtain the deterministic equivalent for the DOF stated in Equation (3.1), which completes the proof of Proposition 3.1. ∎
@@
Let us consider the following decomposition of the predictive risk. Since $\widetilde{X}$ and $\widetilde{\varepsilon}$ are independent from $X_{n}$, we can pursue the following computations to get a new expression for $\hat{r}_{\lambda}(X_{n})$ :
$\hat{r}_{\lambda}(X_{n})$ $=$ $\mathbb{E}[(\widetilde{X}^{t}\hat{\theta}_{\lambda}-\widetilde{X}^{t}\theta_{\ast}-\tilde{\varepsilon})^{2}|X_{n}]$
$=$ $\mathbb{E}[\mathrm{Tr}[\hat{\theta}_{\lambda}\hat{\theta}_{\lambda}^{t}\widetilde{X}\widetilde{X}^{t}+\theta_{\ast}\theta_{\ast}^{t}\widetilde{X}\widetilde{X}^{t}+\tilde{\varepsilon}^{2}-2\theta_{\ast}\hat{\theta}_{\lambda}^{t}\widetilde{X}\widetilde{X}^{t}-2\tilde{\varepsilon}\hat{\theta}_{\lambda}^{t}\widetilde{X}+2\tilde{\varepsilon}\theta_{\ast}^{t}\widetilde{X}]|X_{n}]$
$=$ $\sigma^{2}+\mathrm{Tr}[\mathbb{E}[(\hat{\theta}_{\lambda}\hat{\theta}_{\lambda}^{t}+\theta_{\ast}\theta_{\ast}^{t}-2\theta_{\ast}\hat{\theta}_{\lambda}^{t})|X_{n}]\widetilde{S}_{p}]$
$=$ $\sigma^{2}+\mathbb{E}[\mathrm{Tr}[(\hat{\theta}_{\lambda}-\theta_{\ast})(\hat{\theta}_{\lambda}-\theta_{\ast})^{t}\widetilde{S}_{p}]|X_{n}]$
$=$ $\sigma^{2}+\mathbb{E}[(\hat{\theta}_{\lambda}-\theta_{\ast})^{t}\widetilde{S}_{p}(\hat{\theta}_{\lambda}-\theta_{\ast})|X_{n}]=\sigma^{2}+\hat{R}(\hat{\theta}_{\lambda},\theta_{\ast}),$
with $\hat{R}(\hat{\theta}_{\lambda},\theta_{\ast})=\mathbb{E}[(\hat{\theta}_{\lambda}-\theta_{\ast})^{t}\widetilde{S}_{p}(\hat{\theta}_{\lambda}-\theta_{\ast})|X_{n}]$. Then by using (1.1) and (1.2), we get
$\hat{\theta}_{\lambda}-\theta_{\ast}=-\lambda(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}\theta_{\ast}+(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}\frac{X_{n}^{t}\varepsilon_{n}}{n}.$

### v3

@@ 3.3 Deterministic equivalents of the training and predictive risks
Under Assumptions 1.1 to 1.3, as $\lambda\rightarrow+\infty$, the limit of the deterministic equivalent $r_{\lambda}^{test}(X_{n})$ of the predictive risk is
$\lim_{\lambda\rightarrow+\infty}r_{\lambda}^{test}(X_{n})=\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}.$
Moreover, under Assumptions 1.1 to 1.5, as $\lambda\rightarrow 0$, the limit of the deterministic equivalent $r_{\lambda}^{test}(X_{n})$ is as follows
@@
$\lim_{\lambda\rightarrow 0^{+}}r_{\lambda}^{test}(X_{n})=\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}T_{p}(0^{-})]+\sigma^{2},$
@@
$\lim_{\lambda\rightarrow 0^{+}}r_{\lambda}^{test}$ $=$ $\frac{\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}\kappa(0^{+})(\Sigma_{n}+\kappa(0^{+}))^{-1}]$
$+\frac{\sigma^{2}}{n}\mathrm{Tr}[\kappa^{\prime}(0^{+})\widetilde{S}_{p}\Sigma_{n}(\Sigma_{n}+\kappa(0^{+}))^{-2}]+\sigma^{2},$
where $\kappa^{\prime}(\lambda)$ denotes the derivative with respect to $\lambda$ of the diagonal matrix $\kappa(\lambda)$ defined in Equation (3.2).
In the case $\lambda\rightarrow 0$, Corollary 3.2 yields a deterministic equivalent of the predicitve risk of the mininum norm least-square estimator $\hat{\theta}$. Then, it can be seen that the risk $r_{0}(X_{n})$ exhibit two different behaviors depending on the value of the ratio $p/n$ with respect to one.
If $\frac{p}{n}<1$, then $\hat{\theta}_{0}$ is equivalent to the ordinary least-square estimator $\hat{\theta}=(X_{n}^{\top}X_{n})^{-1}X_{n}^{\top}Y_{n}$ which is known to be unbiased, and the risk $r_{0}(X_{n})$ is thus only composed of a variance term. Moreover, if the variance profile is assumed to be quasi-bistochastic, then, given that
@@
$r_{0}(X_{n})=\frac{\sigma^{2}}{1-\frac{p}{n}}\times\frac{1}{n}\mathrm{Tr}[\widetilde{S}_{p}]+\sigma^{2}=\sigma^{2}\frac{\frac{p}{n}}{1-\frac{p}{n}}+\sigma^{2}\rightarrow\sigma^{2}\frac{c}{1-c}+\sigma^{2},$
as $n\to\infty,\;\frac{p}{n}\to c$, which corresponds, when $c<1$, to the known asymptotic limit of the predictive risk of least squares estimation for iid data when the entries of the features matrix $X_{n}$ are iid centered random variables with variances equal to $1$, see e.g. [HMRT22][Proposition 2].
@@
$r_{0}(X_{n})=\alpha^{2}\left(1-\frac{n}{p}\right)+\sigma^{2}\frac{1}{p/n-1}\rightarrow\alpha^{2}\left(1-\frac{1}{c}\right)+\sigma^{2}\frac{1}{c-1},$
as $n\to\infty,\;\frac{p}{n}\to c$, which corresponds, when $c>1$, to known results on the bias-variance decomposition of the asymptotic limit of the predictive risk of the minimum norm least squares estimator for iid data when the entries of $X_{n}$ are iid centered random variables with variance $1$, see e.g. [HMRT22][Theorem 1].
Beyond the assumption of a quasi-stochastic variance profile, it is difficult to analytically determine the shape of the predictive risk $r_{0}(X_{n})$ as a function of the ratio $p/n$. Indeed, this requires to at least know upper and lower bounds on the magnitude of the diagonal elements of $T_{p}(0^{-})$ and $\kappa(0^{+})$ (and its derivative). As shown by Lemma 3.2 when $p/n<1$, this issue amounts to finding upper and lower bounds of the support of the matrix-valued measure $\mu$ satisfying $T_{p}(-0)=\int_{\tau}^{+\infty}\frac{\mu(dw)}{w}.$ Hence, upper and lower bounding $T_{p}(-0)$ is related to understanding the value of the constant $\tau$ and the size of the support of the limiting spectral distribution of the covariance matrice $\Sigma_{n}$ which remains (to the best of our knowledge) an open problem for random matrices with an arbitrary variance profile.
@@
Under Assumptions 1.1 to 1.3, the function $\lambda\mapsto\hat{r}_{\lambda}^{test}(X_{n})$ reaches its minimium at $\lambda_{\ast}=\frac{\sigma^{2}p}{\alpha^{2}n}$ meaning that $\hat{r}_{\lambda_{\ast}}^{test}(X_{n})\leq\hat{r}_{\lambda}^{test}(X_{n})$, for any $\lambda>0$.
Note that although $\hat{r}_{\lambda}^{test}(X_{n})$ depends on the variance profile, this is not the case of the optimal value of $\lambda_{\ast}$ that minimizes the predictive risk as shown by Corollary 3.3. Moreover, the optimal value $\lambda_{\ast}$ is also the one minimizing $\lambda\mapsto\hat{r}_{\lambda}^{test}(X_{n})$ in the framework of [DW18] with a matrix $X_{n}$ of features made of iid rows.
@@ 4 Numerical experiments
In this section, we illustrate the results of this paper with numerical experiments. Although we do not have an explicit formula for $T_{p}(z)$, this matrix-valued function satisfies the fixed-point equation (2.3). This allows us to approximate $T_{p}(z)$ numerically using a fixed-point algoritheorem. In this manner, we are thus able to obtain a numerical approximation of the deterministic equivalent $r_{\lambda}^{test}(X_{n})$ of the predictive risk. In this section, we consider several variance profiles that we normalize such that
@@
$\Gamma_{n}=\left[\begin{array}[]{ccc}\gamma_{1}^{2}\mathbbm{1}_{n/4}\mathbbm{1}_{p/4}^{\top}&\gamma_{2}^{2}\mathbbm{1}_{n/4}\mathbbm{1}_{p/3}^{\top}&\mathbbm{1}_{n/4}\mathbbm{1}_{5p/12}^{\top}\\ \gamma_{2}^{2}\mathbbm{1}_{n/3}\mathbbm{1}_{p/4}^{\top}&\gamma_{1}^{2}\mathbbm{1}_{n/3}\mathbbm{1}_{p/3}^{\top}&\gamma_{3}^{2}\mathbbm{1}_{n/3}\mathbbm{1}_{5p/12}^{\top}\\ \mathbbm{1}_{5n/12}\mathbbm{1}_{p/4}^{\top}&\gamma_{3}^{2}\mathbbm{1}_{5n/12}\mathbbm{1}_{p/3}^{\top}&\gamma_{1}^{2}\mathbbm{1}_{5n/12}\mathbbm{1}_{5p/12}^{\top}\end{array}\right],$
@@
The values of the predictive and training risks and their deterministic equivalent are compared for various variance profiles in Figures 3 and 3 with $\lambda$ ranging from 0.1 to 10, $n=400$ and $p=600$. The curves displayed in these figures confirm that $r_{\lambda}^{test}(X_{n})$ and $r_{\lambda}^{train}(X_{n})$ are very accurate estimators of $\hat{r}_{\lambda}^{test}(X_{n})$ and $\hat{r}_{\lambda}^{train}(X_{n})$ in high-dimension since the dashed and solid lines coincide. These variance profiles also provide curves of predictive risks having similar shapes (up to a vertical translation). The curves representing the cases of constant and quasi doubly stochastic variance profile coincide, confirming the comment made in Section 1.1 on the doubly stochastic profile and its similarities with the setting of iid data. Moreover, the minimum of $r_{\lambda}^{test}(X_{n})$ is indeed reached at the optimal value $\lambda_{\ast}$ for any variance profile as shown by Corollary 3.3.
Figure 3: Training and predictive risk for several variance profiles with $\lambda$ ranging from $0.1$ to $10$, $\alpha=1$, $\sigma=1$, $n=400$ and $p=600$. (a) Comparison of $\hat{r}_{\lambda}^{test}(X_{n})$ and $r_{\lambda}^{test}(X_{n})$ for several variance profiles. (b) Comparison of $\hat{r}_{\lambda}^{train}(X_{n})$ and $r_{\lambda}^{train}(X_{n})$ for several variance profiles. The dashed lines correspond to the risks while the solid lines correspond to the deterministic equivalents.
We illustrate the appearance of a double or triple descent phenomenon for several variance profiles in Figure 4 (a), (c) and (e). These figures represent the predictive risk and its approximation by a deterministic equivalent for various values of the ratio $p/n$ with $\lambda=0$. The solid lines depicts $r_{0}^{test}$ whereas the dashed lines represent $\hat{r}_{0}^{test}$. For every variance profile, the solid and dashed lines coincide which confirms that $r_{0}^{test}$ is a relevant deterministic equivalent of $\hat{r}_{0}^{test}$. For the constant and quasi doubly stochastic variance profiles, we observe the well known double descent phenomenon as illustrated by Figure 4 (a) since the curves are increasing for $p/n<1$ and decreasing for $p/n>1$. This is related to Corollary 3.2 which states that the expression of $\lim_{\lambda\rightarrow 0}r_{\lambda}^{test}(X_{n})$ depends upon the value of the ratio $p/n$ with respect to one. Nevertheless, for other variance profiles, the shape of the predictive risk can be very different from one variance profile to another, and it differs from the usual double descent. For some variance profiles a phenomenon of triple descent arises, notably for the piecewise constant and block cases, as shown in Figure 4 (c) and (e). We also remark the appearance of a quadruple descent in the case of the polynomial profile, as shown in Figure 4 (e).
@@
Figure 5: Study of MNIST dataset whenever $\lambda=0$. (a) Comparison of real and synthetic data with $\alpha=\sigma=1$, $p=784$ and $n$ varying from $78$ to $1568$. The solid lines correspond to the predictive risk while the dashed line corresponds to the deterministic equivalent.(b) Heatmap of the variance of pixels for each class. These heatmaps serve as variance profiles in the mixture model described below.
We also conducted numerical experiments that highlight the connection between variance-profiled data and mixture models. These experiments are based on the MNIST dataset, a well-known image classification dataset comprising handwritten digits from 0 to 9. This dataset can indeed be viewed as a 10-class mixture model, as described in the introduction, where the $i$-th class corresponds to the digit $i$, and the probability $\pi_{i}$ represents the proportion of digit $i$ in the dataset. Each class exhibits a distinct variance profile, as illustrated in Figure 5 (b). In this figure, we present heatmaps for each class that represent the variance of each pixel within the respective digit images. These heatmaps vary significantly from one class to another, which motivates the use of the variance profile approach. By vectorizing these heatmaps, we extract the variance profiles $S_{i}$. Using these profiles, we then examine the regression model of interest through two distinct approaches. In the first approach, we use real data, where the images from the MNIST dataset serve as feature vectors. Since the images are matrices of size $28\times 28$, we vectorize them into vectors of size 784. In the second approach, we rely on synthetic data by generating random vectors that follow a variance profile $S_{i}$ among those described by the heatmaps in Figure 5 (b). For both scenarios, we consider the case where $\lambda=0$ and compare their predictive risks with the deterministic equivalent derived in our theoretical framework. The results of this comparison are summarized in Figure 5(a). Notably, the curve corresponding to the deterministic equivalent closely matches the one for synthetic data, demonstrating the accuracy of our theoretical predictions. However, these two curves deviate significantly from the one associated with real data. This discrepancy arises because the entries of the feature vectors from real data correspond to the pixels of an image, which are inherently correlated. This violates the assumption of independence among vector entries, a fundamental premise of our study. This limitation underscores one of the primary constraints of our work: we assume independence among the entries of feature vectors, which may be restrictive. Nevertheless, we remain optimistic that our findings can be generalized to scenarios involving correlated data. This optimism is supported by research such as [WCDS12, BGC16], which propose deterministic equivalents for the resolvent matrix in the context of variance-profiled matrices for correlated Gaussian data. Extending our work in this direction could significantly enhance its applicability to real-world datasets.
@@ 5 Conclusion
In this paper, we derive a deterministic equivalent of the DOF and the predictive risk of ridge (less) regression in a high-dimensional framework with a variance profile to handle the setting of non-iid data. We use RMT to determine a deterministic equivalent of the diagonal of the resolvent matrix $Q_{p}(z)$. Our work extends the study of the DOF and the predictive risk from [DW18] and [Bac24] to the case of general variance profile. It appears that also the result of [DW18] and [Bac24] still hold in the case of a quasi doubly stochastic profile. However, there are variance profiles that cause a behavior of the predictive risk of the minimum norm least square estimator that differs from the double descent phenomenon as classically observed in the case of a constant variance profile. Our numerical experiments confirm that our deterministic equivalent accurately estimates the predictive risk in high-dimension. Our results also allow to understand how assuming such a variance profile for the data influences the statistical properties of ridge regression when compared to the standard assumption of iid observations. We hope that our approach on the use of variance profiles may lead to further research works on the statistical analysis of other estimators than ridge regression in more complex models with non-iid data.
@@
A.1 Proofs of Proposition 3.1 and Lemmas 3.1 and 3.1
@@
Thanks to Proposition 2.1, one has that $\frac{1}{p}\widehat{df}_{1}(\lambda)\sim 1-\lambda g_{\nu_{n}}(-\lambda)$. Let us now remark that, by Proposition 2.1, one has that
@@
where $\kappa_{j}(\lambda)=\frac{\operatorname{Tr}[\widetilde{D}_{j}]}{\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(-\lambda)]}$, for $1\leq j\leq p,$ and $\kappa(\lambda)=\mathop{\rm diag}\nolimits(\kappa_{1}(\lambda),\ldots,\kappa_{p}(\lambda)).$
Consequently, we obtain the deterministic equivalent for the DOF stated in Equation (3.1), which completes the proof of Proposition 3.1. ∎
@@
Let’s compute a new expression for $\hat{r}_{\lambda}^{train}(X_{n})$
$\hat{r}_{\lambda}^{train}(X_{n})$ $=$ $\frac{1}{n}\mathbb{E}[\parallel Y_{n}-X_{n}^{\top}\hat{\theta}_{\lambda}\parallel^{2}|X_{n}]$
$=$ $\frac{1}{n}\mathbb{E}[\parallel(I_{n}-\frac{1}{n}X_{n}^{\top}X_{n}Q_{p}(-\lambda))Y_{n}\parallel^{2}|X_{n}]$
$=$ $\frac{\lambda^{2}}{n}\mathbb{E}[\parallel Q_{p}(-\lambda)Y_{n}\parallel^{2}|X_{n}]$
$=$ $\frac{\lambda^{2}}{n}\mathbb{E}[\operatorname{Tr}[Q_{p}(-\lambda)X\beta_{\ast}\beta_{\ast}^{\top}X_{n}^{\top}Q_{p}(-\lambda)+2Q_{p}(-\lambda)X_{n}\beta_{\ast}\varepsilon^{\top}Q_{p}(-\lambda)$
$+Q_{p}(-\lambda)\varepsilon\varepsilon^{\top}Q_{p}(-\lambda)]|X_{n}]$
$=$ $\frac{\lambda^{2}\alpha^{2}}{n}\mathrm{Tr}\Bigg{[}Q^{2}_{p}(-\lambda)\frac{X_{n}X_{n}^{\top}}{p}\Bigg{]}+\frac{\lambda^{2}\sigma^{2}}{n}\mathrm{Tr}[Q^{2}_{p}(-\lambda)]$
$=$ $\frac{\lambda^{2}\alpha^{2}}{n}\mathrm{Tr}\Bigg{[}Q_{p}(-\lambda)-\lambda Q_{p}^{\prime}(-\lambda)\Bigg{]}+\frac{\lambda^{2}\sigma^{2}}{n}\mathrm{Tr}[Q_{p}^{\prime}(-\lambda)].$
Let us consider the following decomposition of the predictive risk. Since $\tilde{x}$ and $\widetilde{\varepsilon}$ are independent from $X_{n}$, we can pursue the following computations to get a new expression for $\hat{r}_{\lambda}^{test}(X_{n})$:
$\hat{r}_{\lambda}^{test}(X_{n})$ $=$ $\mathbb{E}[(\tilde{x}^{\top}\hat{\theta}_{\lambda}-\tilde{x}^{\top}\beta_{\ast}-\tilde{\varepsilon})^{2}|X_{n}]$
$=$ $\mathbb{E}[\mathrm{Tr}[\hat{\theta}_{\lambda}\hat{\theta}_{\lambda}^{\top}\tilde{x}\tilde{x}^{\top}+\beta_{\ast}\beta_{\ast}^{\top}\tilde{x}\tilde{x}^{\top}+\tilde{\varepsilon}^{2}-2\beta_{\ast}\hat{\theta}_{\lambda}^{\top}\tilde{x}\tilde{x}^{\top}-2\tilde{\varepsilon}\hat{\theta}_{\lambda}^{\top}\tilde{x}$
$+2\tilde{\varepsilon}\beta_{\ast}^{\top}\tilde{x}]|X_{n}]$
$=$ $\sigma^{2}+\mathrm{Tr}[\mathbb{E}[(\hat{\theta}_{\lambda}\hat{\theta}_{\lambda}^{\top}+\beta_{\ast}\beta_{\ast}^{\top}-2\beta_{\ast}\hat{\theta}_{\lambda}^{\top})|X_{n}]\widetilde{S}_{p}]$
$=$ $\sigma^{2}+\mathbb{E}[\mathrm{Tr}[(\hat{\theta}_{\lambda}-\beta_{\ast})(\hat{\theta}_{\lambda}-\beta_{\ast})^{\top}\widetilde{S}_{p}]|X_{n}]$
$=$ $\sigma^{2}+\mathbb{E}[(\hat{\theta}_{\lambda}-\beta_{\ast})^{\top}\widetilde{S}_{p}(\hat{\theta}_{\lambda}-\beta_{\ast})|X_{n}]=\sigma^{2}+\hat{R}(\hat{\theta}_{\lambda},\beta_{\ast}),$
with $\hat{R}(\hat{\theta}_{\lambda},\beta_{\ast})=\mathbb{E}[(\hat{\theta}_{\lambda}-\beta_{\ast})^{\top}\widetilde{S}_{p}(\hat{\theta}_{\lambda}-\beta_{\ast})|X_{n}]$. Then by using (1.1) and (1.2), we get
$\hat{\theta}_{\lambda}-\beta_{\ast}=-\lambda(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}\beta_{\ast}+(\widehat{\Sigma}_{n}+\lambda I_{p})^{-1}\frac{X_{n}^{\top}\varepsilon_{n}}{n}.$